OpenStudy (anonymous):

The vertex of the parabola below is at the point (1, -3). Which of the equations below could be the one for this parabola?

OpenStudy (anonymous):

A. y = (x + 3)2 + 1 B. x = -3(y - 1)2 C. y = (x - 1)2 - 3 D. y = (x + 1)2 - 3

OpenStudy (math&ing001):

Plug x and y with their values 1 and -3 respectively in the equations, and see which one checks.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@math&ing001 is correct just plug in.

OpenStudy (anonymous):

A?

OpenStudy (loser66):

@oscar55773 why don't you "plug in" like what the helper suggested? why don't you stop guessing?

OpenStudy (anonymous):

A is wrong by a long shot

OpenStudy (loser66):

Or you don't know what the "plug in" is?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

(X,Y) = (1,-3)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

y=(x,h)+k

OpenStudy (math&ing001):

Here's an example: y =(x + 1)^2 - 3 -3 =(1 + 1)^2 - 3 -3 =2^2 - 3 -3 = 1 So D is not an option. Do the same for the others.